The length of a rectangle is 6in longer than its width.
If the perimeter of the rectangle is 44in , find its area.

Answer by Narsinghchauhan8960103366(17)

Let the width of rectangle is x in
So the length is (x+6) in
According to question
Perimeter of Rectangle =2(length+width)
44=2[(x+6)+x]
44=2(2x+6)
44=4x+12
4x=44-12=32
x=32/4=8 in
width =8 in
length =8+6=14
Perimeter of Rectangle = length*width
=14*8=112 square in